### v1

# CardioLab: Laboratory Values Estimation from Electrocardiogram Features - An Exploratory Study

Abstract
Introduction: Laboratory value represents a cornerstone of medical diagnostics, but suffers from slow turnaround times, and high costs and only provides information about a single point in time. The continuous estimation of laboratory values from non-invasive data such as electrocardiogram (ECG) would therefore mark a significant frontier in healthcare monitoring. Despite its transformative potential, this domain remains relatively underexplored within the medical community. Methods: In this preliminary study, we used a publicly available dataset (MIMIC-IV-ECG) to investigate the feasibility of inferring laboratory values from ECG features and patient demographics using tree-based models (XGBoost). We define the prediction task as a binary prediction problem of predicting whether the lab value falls into low or high abnormalities. The model performance can then be assessed using AUROC. Results: Our findings demonstrate promising results in the estimation of laboratory values related to different organ systems based on a small yet comprehensive set of features. While further research and validation are warranted to fully assess the clinical utility and generalizability of ECG-based estimation in healthcare monitoring, our findings lay the groundwork for future investigations into approaches to laboratory value estimation using ECG data. Such advancements hold promise for revolutionizing predictive healthcare applications, offering faster, non-invasive, and more affordable means of patient monitoring.
1 Introduction
Abnormal laboratory values play a crucial role in clinical settings as they indicate underlying health conditions, sometimes posing severe and potentially life-threatening risks, especially in cases of serious illnesses [1, 2]. These abnormal values significantly increase the risk of mortality compared to normal ranges [3], highlighting the need for vigilant monitoring and effective management strategies. Continuous monitoring of patients often involves inherent limitations such as invasive procedures like frequent blood sampling, which are resource-intensive [4] and suffer from delays between sampling and obtaining results [5] or potentially missing staffing of laboratory medicine during the night, thereby limiting real-time monitoring capabilities.
While previous research has explored the correlation between abnormal laboratory values and changes in electrocardiogram (ECG) readings [6, 7], the full potential of utilizing ECG data for accurately estimating these abnormalities remains largely untapped. Therefore, in this work, we propose a first exploratory study in which we investigate the feasibility of estimating laboratory values from ECG features with the addition of other data modalities such as patient demographics.
2 Background
Significance of abnormal laboratory values In the United States, annual costs for preoperative laboratory testing alone amount to \$18 billion [4]. In intensive care units (ICUs), these costs can rise to \$14 billion, constituting up to 25% of total ICU expenditures [8]. Identifying abnormal lab values is crucial for patient care. In elderly surgical patients, preoperative tests often show abnormal creatinine (12%), hemoglobin (10%), and glucose (7%) levels [3]. In ICUs, these tests are essential for adjusting drug dosages and procedures for critically ill patients, playing a key role in treatment strategies [5].
Monitoring and management strategies Current guidelines for monitoring patients with abnormal lab values recommend instruments and methods for early detection, continuous monitoring, and effective interventions. They include organ-specific tests, like liver function [9], and are adapted for settings such as the emergency department (ED) [10]. While traditional phlebotomy is common, non-invasive alternatives are emerging for biomarkers like glucose [11] and hemoglobin [12].
Integration of ECG data for values prediction The use of ECG data for laboratory values estimation has been briefly investigated, where previous works demonstrated predictive capabilities in glucose [13], serum potassium [14], and electrolyte imbalances [15, 16]. Lastly, other predictive modalities such as vital signs from wearable sensors were investigated to predict laboratory values [17].
3 Methods
Dataset The dataset comprises data sourced from the publicly available MIMIC-IV [18, 19] and MIMIC-IV-ECG [20, 19] datasets. It includes a comprehensive set of non-invasive features: demographics such as gender, age, and race (Caucasian, African, Asian, Latino, or other); ECG features including RR interval, P onset, P end, QRS onset, QRS end, and T end (all in milliseconds), as well as P axis, QRS axis, and T axis (in degrees). In terms of target abnormalities, we work on binary classification cases where we define a positive case when the considered laboratory value is lower or higher than the median low or high threshold values (provided as part of MIMIC-IV) across all samples of the same laboratory value. For dataset splitting, we follow the stratification based on patients, age, and gender from the previously proposed MIMIC-IV-ECG-ICD [21] splits with a train, validation, and test ratio of 18:1:1 which ensures balanced diagnoses in each of the folds. Finally, for this work, we consider final laboratory value cases where we obtain at least 10 positive and 10 negative cases per fold.
Models and performance evaluation We fit and train individual extreme gradient boosting (XGBoost) tree models with a max depth of 1 per laboratory value in a binary classification setting. We evaluate performances on the macro average across all areas under the respective receiver operating curves (AUROC) (macro AUROC), where to assess statistical uncertainty resulting from the finite size and specific composition of the test set, we employ empirical bootstrap on the test set with $n=1000$ iterations and report 95% confidence intervals.
4 Results
Table 1: Performance results for laboratory values with AUROC > 0.70, presented by value name, abnormality setting, threshold defining abnormality by the number of units, (units), number of samples, class prevalence, and AUROC with 95% confidence intervals. All of these values represent blood fluid.
Value Threshold Unit Samples [Prev.] AUROC (95% CI)
Urea Nitrogen <6.0 mg/dL 240892 [2139] 0.856 (0.855, 0.857)
NTproBNP >353.0 pg/mL 23206 [18035] 0.848 (0.848, 0.849)
Hematocrit >51.0 % 238567 [1381] 0.775 (0.773, 0.777)
Cortisol >20.0 ug/dL 982 [486] 0.771 (0.772, 0.781)
Osmolality <275.0 mOsm/kg 2784 [730] 0.765 (0.763, 0.768)
Hemoglobin >17.5 g/dL 236802 [802] 0.754 (0.753, 0.759)
Creatinine <0.5 mg/dL 241968 [2499] 0.742 (0.741, 0.744)
Urea Nitrogen >20.0 mg/dL 240892 [100216] 0.736 (0.736, 0.736)
Creatinine >1.2 mg/dL 241968 [69012] 0.735 (0.735, 0.736)
PT <9.4 sec 137917 [380] 0.732 (0.728, 0.734)
C-Reactive Protein >5.0 mg/L 3520 [2107] 0.730 (0.728, 0.734)
Anion Gap >20.0 mEq/L 230666 [18299] 0.726 (0.725, 0.727)
Cholesterol, HDL <41.0 mg/dL 9023 [2103] 0.725 (0.725, 0.729)
Lymphocytes >42.0 % 171441 [9591] 0.722 (0.721, 0.723)
RDW-SD >46.3 fL 73001 [35887] 0.720 (0.719, 0.721)
Bands >5.0 % 16197 [2587] 0.719 (0.718, 0.721)
Bicarbonate, Whole Blood <21.0 mEq/L 8039 [1720] 0.715 (0.712, 0.717)
Ferritin <13.0 ng/mL 4643 [438] 0.714 (0.714, 0.722)
Lymphocytes <18.0 % 171441 [73323] 0.713 (0.712, 0.713)
Monocytes >11.0 % 9898/171442 0.713 (0.712, 0.714)
Bilirubin, Direct > 0.3 mg/dL 3438 [1968] 0.708 (0.704, 0.710)
Albumin <3.5 g/dL 57040 [12834] 0.708 (0.707, 0.708)
Sodium >145.0 mEq/L 235578 [5361] 0.706 (0.705, 0.707)
eAG >123.0 mg/dL 9485 [4916] 0.704 (0.702, 0.705)
Ferritin >150.0 ng/mL 4643 [2013] 0.704 (0.701, 0.705)
% Hemoglobin A1c >5.9 % 10665 [5651] 0.700 (0.699, 0.701)
Overall Predictive Performance Table 1 contains the performance results in terms of AUROC for individual laboratory value abnormalities above 0.70. This includes the threshold for their label definition, units, number of samples, and prevalence. Overall, low urea nitrogen and NTproBNP, achieve an AUROC above 0.80. The table lists 26 laboratory values with AUROC greater than 0.70, indicating strong predictive performance. These values span various physiological systems, demonstrating the model’s capability to predict abnormalities across diverse domains beyond cardiovascular (NTproBNP, hematocrit, hemoglobin), coagulation (PT), and electrolyte balance (osmolality, anion gap, sodium). Notably, the model also encompasses renal (urea nitrogen, creatinine), endocrine (cortisol, eAG), immune (bands, monocytes, lymphocytes), inflammatory response (C-Reactive protein), metabolic (cholesterol, bicarbonate), and respiratory (pO2) systems. This broad applicability underscores the model’s extensive utility in clinical prediction across a wide range of physiological parameters.
Low urea nitrogen use-case: predictive performance for different patient demographics By gender, males perform better than females despite having a smaller prevalence rate. By race, the groups of Caucasians and others perform above the overall population mean, Caucasians however, with the largest prevalence, Asians have the lowest predictive performance, however, also the ones with the fewest samples, to make a fair comparison, we down-sampled the Caucasian training population to Asians population, and conclude that indeed Caucasian perform better than (at least) Asians, even with the same number of samples. By age, the 65 to 77 years age group is the most predictive across all groups despite having a small prevalence.
Low urea nitrogen use-case: interpretability XGBoost enables explainable methods like Shapley additive explanations (SHAP) [22]. Using SHAP, we identify ECG features contributing most to abnormality predictions. Literature indicates hemodialysis patients with reduced urea nitrogen levels often show ECG abnormalities [23]. Our findings suggest: (1) High T end values (long T wave) correlate with prolonged repolarization abnormalities, potentially linked to observed ST segment changes (depression or elevation) and inverted T waves in these patients. (2) Low T axis values correlate with T wave inversions and other repolarization abnormalities. (3) High QRS axis values indicate a shift in ventricular depolarization’s mean electrical axis, possibly associated with ventricular hypertrophy or conduction abnormalities in hemodialysis patients. Furthermore, high NTproBNP values, our second-best performer, align with literature associating abnormal ECG features in a specific cohort [24].
5 Discussion
Clinical significance This work enhances ECG predictive capabilities beyond traditional applications, laying the groundwork for comprehensive patient monitoring as in [25] for diagnostics. By analyzing ECG data alongside patient demographics, we predict abnormalities in crucial laboratory markers such as electrolyte levels (e.g., Anion gap, Sodium, Bicarbonate) and renal function indicators like urea nitrogen and creatinine. This predictive approach facilitates early detection of conditions such as arrhythmias due to electrolyte imbalances or kidney dysfunction, enabling proactive interventions that can optimize patient care by providing faster and low-cost results [26].
Limitations We identify several limitations in our current work, firstly, for clinical consideration, one might need to externally validate these models on a separate population cohort. Secondly, we aim to estimate the closest absolute value within a horizon threshold, however, fixed intervals between features and targets for all samples might mitigate variance. Lastly, A more fine-grained setting can include drug-specific pharmacokinetic information to account for drug effects in the target space. However, the complexity of it would increase significantly as half-life, and mechanisms of action and elimination are drug-specific.
Future work While this work investigates the predictive performance of laboratory values by discriminating ECG features within a stratified population mostly presenting diagnoses [21], it is also important to understand how changes in ECG features contribute to predictions in a population of only healthy individuals [27]. Additionally, previous research demonstrates the superiority of using ECG signals over features for diagnosis and decompensation prediction [28]. Extending this work to include waveforms rather than relying solely on tabular data would be a valuable direction for future research.
Data and code availability Code for dataset preprocessing and experimental replications can be found in our dedicated repository [29].
Appendix
Table 2: Performance results by demographic groups, presented by category, the number of samples, prevalence, and AUROC with 95% confidence intervals.
Category Samples [Prev.] AUROC
Gender
Males 122,125 [686] 0.844 (0.842, 0.846)
Females 118,767 [1453] 0.829 (0.828, 0.831)
Race
Caucasians 157,926 [1287] 0.881 (0.880, 0.882)
Africans 40,205 [432] 0.784 (0.784, 0.790)
Asians 7,295 [63] 0.653 (0.652, 0.666)
Latinos 13,842 [149] 0.763 (0.765, 0.774)
Other 17,306 [206] 0.871 (0.869, 0.873)
Age by quantiles
(18-49yo.) 54,830 [1181] 0.723 (0.720, 0.723)
(50-64yo.) 54,203 [617] 0.702 (0.701, 0.708)
(65-77yo.) 65,854 [227] 0.808 (0.805, 0.815)
(>=78yo.) 52,596 [114] 0.709, (0.705, 0.708)